# Online Fair Allocations with Binary Valuations and Beyond

### v1

@@ 10                  $\pi=(\pi_{1},\cdots,\pi_{i-1},\pi_{i+1},\cdots,\pi_{n},\pi_{i})$;
For submodular binary valuation functions, no deterministic online algorithm can achieve $(\frac{1}{2}+\epsilon)$-max-USW and non-wastefulness for any $\epsilon>0$.
Proof.
Consider an online allocation instance with two agents $\{1,2\}$. For the first arrived item $e_{1}$, the marginal utilities for agents $1$ and $2$ are both $1$, $v_{1}(\{e_{1}\})=v_{2}(\{e_{1}\})=1$. Without loss of generality, we allocate $e_{1}$ to agent $1$. When the second item $e_{2}$ arrives, $v_{1}(\{e_{2}\})=v_{1}(\{e_{1},e_{2}\})=1$,$v_{2}(\{e_{2}\})=0$, the marginal utilities for agents $1$ and $2$ are both 00. No matter which agent gets $e_{2}$, $USW(A^{2})=1$. However, the max-USW is $2$ by allocating $e_{1}$ to agent $2$ and $e_{1}$ to agent $1$. ∎
There is a side result of our algorithm, where the Marginal-Greedy Algorithm (Algorithm 2) can also work for additive binary valuation functions to compute an NW allocation that is EF1, MMS, and max-USW, where for each agent $i\in N$, $v_{i}$ is additive and $v_{i}(e)\in\{0,1\}$ for any $e\in T$.
3.3 Additive Bi-valued Valuations
In this part, we consider additive bi-valued valuation functions. For any agent $i\in N$, $v_{i}$ is bi-valued if $v_{i}(e)\in\{a_{i},b_{i}\}$ for any $e\in T$, where $0<a_{i}\leq b_{i}$. Someone may wonder whether a non-wasteful EF1 or MMS allocation can be found with additive bi-valued valuation functions because the additive binary valuation function can be seen as a special additive bi-valued valuation function. However, we find that even for two agents with additive bi-valued valuations, a non-wasteful EF1 or MMS allocation cannot be guaranteed.
3.3.1 Two Agents
We study the setting of additive bi-valued valuations for two agents and establish the best possible result for the approximation of EF1 and MMS simultaneously. In order to achieve this goal, we propose the Adapted Envy-Graph Procedures (Algorithm 3). Note that there are two technical difficulties for us when designing this approximation algorithm. One is that we do not know the value of future items. The other is that we cannot do any reallocation.
The main idea of our algorithm is that we take the Adapted Envy-Graph Procedures as the base algorithm to allocate items. If we follow the base algorithm, there will be an envy cycle between two agents, and we temporarily terminate it. Then, we consider implementing the Preliminary-Breaking-Cycle Algorithm (Algorithm 4) to eliminate the potential envy cycle manually. Then, there are two possible results of Algorithm 4. One is that we have eliminated the potential envy cycle successfully. The other is that there is still an envy cycle, and we execute the Deep-Breaking-Cycle Algorithm (Algorithm 5) to eliminate it further. In whichever case, when there is no envy cycle and the allocation is EF1, we go back to the base algorithm. The illustration of Algorithms 3, 4 and 5 is shown in Figure 2. Before presenting the theorem, we first define an envy graph.
Figure 2: The illustration of Algorithms 3, 4 and 5.
Definition 6 (Envy Graph for Goods).
Given an allocation $A^{k}$, in an envy graph G, each vertex corresponds to an agent, and each directed edge $(i,j)$ represents that agent $i$ envies agent $j$, where we have $v_{i}(A_{i}^{k})<v_{i}(A_{j}^{k})$.
Input: An instance $(T,N,(v_{1},v_{2}))$ with additive bi-valued valuation functions
Output: An non-wasteful approximate EF1 and approximate MMS allocation $A$
@@ 3.3.1 Two Agents
2Let $A=(\emptyset,\ldots,\emptyset)$ and build an envy graph $G$ for two agents;
3 when item $e\in T$ arrives do
4       if  there is no envy between the two agents  then
5             if $\forall k\in N$, $v_{k}(e)=a_{k}$ or $\forall k\in N$, $v_{k}(e)=b_{k}$  then
6                   Arbitrarily choose one agent $i\in N$;
@@
8                   Choosing the agent $i\in N$ who has the valuation of $b_{i}$;
@@
10
11      else
12             Let $i\in N$ be an unenvied agent and $j=2-i$;
@@
14                   Terminate this algorithm and implement the Preliminary-Breaking-Cycle algorithm (Algorithm 4);
@@
17      $A_{i}=A_{i}\cup\{e\}$;
18
return $A$;
Algorithm 3 Adapted Envy-Graph Procedures
Theorem 5.
For the deterministic allocation of indivisible goods for two agents with additive bi-valued valuations, the Adapted Envy-Graph Procedures (Algorithm 3) computes an NW allocation that satisfies $\frac{1}{2}$-EF1 and $\frac{1}{3}$-MMS.
Before we show Theorem 5, we first give the following key observations and lemmas.
@@ 17      $A_{i}=A_{i}\cup\{e\}$;
When Algorithm 4 is executed, for the first arriving item $e$, we have $v_{i}(e)=a_{i}$.
@@
The reason why Algorithm 4 is executed is that for the arriving item $e$, if we add it to agent $i$’s bundle, there will be an envy cycle between two agents. i.e., $v_{i}(A_{i}\cup\{e\})<v_{i}(A_{j})$ and $v_{j}(A_{j})<v_{i}(A_{i}\cup\{e\})$. Suppose for the contradiction that $v_{i}(e)=b_{i}$. Since the input allocation $A$ is EF1, for agent $i$, we have $v_{i}(A_{j})-v_{i}(A_{i})\leq b_{i}$. Thus, allocating item $e$ to agent $i$ eliminates the envy and cannot lead to an envy cycle, which is a contradiction. Therefore, for the arriving item $e$, it holds that $v_{i}(e)=a_{i}$. ∎
@@
When Algorithm 4 is executed, we have $|A_{i}|\geq 1$ and $|A_{j}|\geq 1$.
@@
Suppose that $|A_{i}|=0$. In that case, we have $|A_{j}|=1$. Otherwise, there would be an envy cycle. After allocating $e$, there will be an envy cycle, i.e., $v_{i}(e)=a_{i}<v_{i}(A_{j})=b_{i}$ and $v_{j}(A_{j})<v_{j}(e)$. If the second inequality holds, we have $v_{j}(A_{j})=a_{j}$ but $v_{i}(A_{j})=b_{i}$, which contradicts the line 3 in Algorithm 3. Suppose that $|A_{j}|=0$. In that case, we have $|A_{i}|=0$, which contradicts the fact that after allocating $e$, an envy cycle appears. ∎
@@
Algorithm 4 computes an NW allocation that is EF1 with or without an envy cycle, or $\frac{1}{2}$-EF1.
@@
Before Algorithm 4 is implemented, the output allocation $A=(A_{1},A_{2})$ by Algorithm 3 is EF1. According to the algorithm description, it is executed in at most two rounds.
@@
$\lambda=1$. For the first arriving item $e_{1}$, by Observation 1, we have $v_{i}(e_{1})=a_{i}$. If $\lvert A_{j}\rvert=1$, we have $v_{i}(A_{j})=b_{i}$ since agent $i$’s bundle is EF1. Then, agent $j$ picks this item. For agent $i$, her bundle is EF1 since we have $\lvert A_{i}\lvert\geq 1$ and $v_{i}(e_{1})=a_{i}$. For agent $j$, her new bundle is EF since she does not envy agent $i$. Thus, this allocation is EF1 without an envy cycle, and the algorithm is terminated. If $\lvert A_{j}\rvert>1$, let us consider two cases about the valuation of item $e_{1}$.
@@
  • If $v_{j}(e_{1})=b_{j}$, item $e_{1}$ is allocated to agent $j$ and we set $\mu=1$. In this case, for agent $j$, her new bundle is EF since $v_{j}(A_{j}\cup\{e_{1}\})>v_{j}(A_{j})\geq v_{j}(A_{i})$ and for agent $i$, her bundle is $\frac{1}{2}$-EF1, i.e, $v_{i}(A_{i})\geq\frac{1}{2}v_{i}(A_{j}\cup\{e_{1}\}\setminus\{g\})$ holds for $g\in A_{j}\cup\{e_{1}\}$, since we have $v_{i}(A_{i})\geq v_{i}(A_{j}\setminus\{g\})$ for some $g\in A_{j}$ and $v_{i}(A_{i})\geq a_{i}=v_{i}(e_{1})$. Thus, this allocation is $\frac{1}{2}$-EF1.
@@
  • If $v_{j}(e_{1})=a_{j}$, item $e_{1}$ is allocated to agent $i$ and we set $\mu=2$. In this case, for agent $i$, her new bundle is still EF1 since she is an unenvied agent in Algorithm 3 and the input allocation $A$ is EF1, and for agent $j$, her bundle is EF1 since we have $v_{j}(A_{j})\geq v_{j}(A_{i})$ but $v_{j}(A_{j})<v_{j}(A_{j}\cup\{e_{1}\})$. Thus, this allocation is EF1 with an envy cycle.
$\lambda=2$. The second item $e_{2}$ arrives. Let us consider two cases about $\mu$.
@@
  • If $\mu=1$, agent $i$ picks it. In this case, for agent $j$, her bundle is still EF since we have $v_{j}(A_{j})\geq v_{j}(A_{i})$ and $v_{j}(e_{1})=b_{j}\geq v_{j}(e_{2})$. For agent $i$, her new bundle is EF1 since we have $v_{i}(A_{i})\geq v_{i}(A_{j}\setminus\{g\})$ for some $g\in A_{j}$ and $v_{i}(e_{2})\geq v_{i}(e_{1})=a_{i}$. Thus, this allocation is EF1 without an envy cycle, and the algorithm is terminated.
@@
  • If $\mu=2$, let us consider two cases. When $v_{i}(e_{2})=a_{i}$, agent $j$ picks it. For agent $i$, her bundle is still EF1 since we have $v_{i}(A_{i})\geq v_{i}(A_{j}\setminus\{g\})$ for some $g\in A_{j}$ and $v_{i}(e_{1})=v_{i}(e_{2})=a_{i}$. For agent $j$, her new bundle is EF because $v_{j}(A_{j})\geq v_{j}(A_{i})$ and $v_{j}(e_{2})=v_{j}(e_{1})=a_{j}$. Hence, the allocation is EF1 without an envy cycle, and the algorithm is terminated. When $v_{i}(e_{2})=b_{i}$, it is allocated to agent $i$. For agent $i$, her new bundle is EF since it holds that $v_{i}(A_{i}\cup\{e_{1},e_{2}\})=v_{i}(A_{i})+a_{i}+b_{i}\geq v_{i}(A_{j}\setminus\{g\})+v_{i}(g)+a_{i}$ for some $g\in A_{j}$ where the inequality follows from the fact that $v_{i}(g)\in\{a_{i},b_{i}\}$. For agent $j$, her bundle is $\frac{1}{2}$-EF1, i.e., $v_{j}(A_{j})\geq\frac{1}{2}v_{j}(A_{i}\cup\{e_{1},e_{2}\}\setminus\{g\})$ holds for some $e\in A_{i}\cup\{e_{1},e_{2}\}$, since $v_{j}(A_{j})\geq v_{j}(A_{i})$ and $v_{j}(A_{j})\geq v_{j}(e_{1})=\min\{v_{j}(e_{1}),v_{j}(e_{2})\}$. Thus, this allocation is $\frac{1}{2}$-EF1 and the algorithm is terminated.
@@
When Algorithm 5 is executed, there are two properties about input allocation $A$:
@@
  • Let $e_{1}$ and $e_{2}$ denote two new allocated items in Algorithm 4, we have $v_{j}(A_{j})\geq\frac{1}{2}v_{j}(A_{i}\setminus\{e\})$, where $e\in\arg\max_{g\in\{e_{1},e_{2}\}}v_{j}(g)$.
@@
When Algorithm 5 is ready to be executed, Algorithm 4 is terminated for $\lambda=\mu=2$. Then, by the proof of Algorithm 4, it is not hard to see that the statements are correct. ∎
@@
Algorithm 5 computes an NW allocation that is EF1 with or without an envy cycle, or $\frac{1}{2}$-EF1.
@@
By Observation 3, the input allocation $A=(A_{1},A_{2})$ is $\frac{1}{2}$-EF1, where $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}$ and $v_{j}(A_{j})\geq\frac{1}{2}v_{j}(A_{i}\setminus\{e\})$ for some $e\in A_{i}$. Unlike Algorithm 4, this algorithm may be implemented in many rounds. Let $g_{1}$ and $g_{2}$ denote the two online items that are allocated to agent $i$ in Algorithm 4 and we have $\min\{v_{j}(g_{1}),v_{j}(g_{2})\}=a_{j}$.
$\lambda=1$. For the arriving item $e_{1}$ when $\lambda=1$, let us consider four cases of agents’ valuations.
  • When $v_{i}(e_{1})=a_{i}$ and $v_{j}(e_{1})=b_{j}$, agent $j$ picks it and we set $\mu=1$. For agent $i$, her bundle is EF since we have $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}=v_{i}(A_{j}\cup\{e_{1}\})$. For agent $j$, her new bundle is EF1 since we have $v_{j}(A_{j})\geq v_{j}(A_{i}\setminus\{g_{1},g_{2}\})$ and $v_{j}(e_{1})=b_{j}\geq\max\{v_{j}(g_{1}),v_{j}(g_{2})\}$. Thus, this allocation is EF1 without an envy cycle.
  • When $v_{i}(e_{1})=b_{i}$ and $v_{j}(e_{1})=b_{j}$, it is allocated to agent $j$ and we set $\mu=2$. For agent $i$, her bundle is EF1 since after removing $e_{1}$ from $A_{j}$, we have $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}$. For agent $j$, her new bundle is also EF1 since we have $v_{j}(A_{j})\geq v_{j}(A_{i}\setminus\{g_{1},g_{2}\})$ and $v_{j}(e_{1})=b_{j}\geq\max\{v_{j}(g_{1}),v_{j}(g_{2})\}$. Thus, this allocation is EF1 with an envy cycle.
  • When $v_{i}(e_{1})=a_{i}$ and $v_{j}(e_{1})=a_{j}$, it is allocated to agent $j$ and we set $\mu=3$. For agent $i$, her bundle is EF since we have $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}=v_{i}(A_{j}\cup\{e_{1}\})$. For agent $j$, her new bundle is EF1 since we have $v_{j}(A_{j})\geq v_{j}(A_{i}\setminus\{g_{1},g_{2}\})$ and $v_{j}(e_{1})=a_{i}=\min\{v_{j}(g_{1}),v_{j}(g_{2})\}$. Thus, this allocation is EF1 without an ency cycle.

### v2

@@ 20  Update the envy graph $G$;
For the deterministic allocation of indivisible goods for two agents with additive personalized bi-valued valuations, the AEGP algorithm (Algorithm 3) computes an NW allocation that satisfies $\frac{1}{2}$-EF1 and $\frac{1}{3}$-MMS.
The following two lemmas are crucial in the proof of Theorem 4.
Input: $N,A,\alpha,e$
Output: An NW allocation A and $\alpha$
@@
2if there is no envy between the two agents  then
3 if $\exists i\in N$, $v_{i}(e)=b_{i}$  then
4      Choosing the agent $i$ who values $e$ at $b_{i}$ (breaking ties arbitrarily);
@@
6      Arbitrarily choose one agent $i$;
@@
8
9else
10   Let $i\in N$ be an unenvied agent and j=2ij=2-i;
@@
12    $\alpha=2$;
13    return A and $\alpha$;
14   end if
@@ 12    $\alpha=2$;
[…]
return A and $\alpha$;
Algorithm 4 EGP (Envy-Graph Procedure)
Lemma 3.
The EGP algorithm (Algorithm 4) computes an NW allocation that is EF1 without an envy cycle.
Proof.
Note that there are three kinds of input allocation. The first one is the initial empty allocation, which is EF1 without an envy cycle. The second one is the output allocation of the PBC algorithm. The last one is the output allocation of the DBC algorithm. It suffices to show that the last two kinds of input allocation are EF1 without an envy cycle. That is because if the input allocation of the EGP algorithm is EF1 without an envy cycle, by the description of the algorithm and induction, it is easy to check that the property of EF1 without an envy cycle is always maintained. By Lemma 4, we can derive that when the PBC or DBC algorithm outputs $\alpha=1$, implying that the EGP algorithm is ready to be executed, the output allocation of the PBC or DBC algorithm is EF1 without an envy cycle. Therefore, our claim holds. ∎
Lemma 4.
The PBC or DBC algorithm (Algorithm 5 or 6) computes an NW allocation satisfying
  • $\frac{1}{2}$-EF1,
  • or EF1 with an envy cycle,
  • or EF1 without an envy cycle.
We split the proof of Lemma 4 into two parts, where we show that the claim holds for both the PBC and DBC algorithms.
@@
When the PBC algorithm (Algorithm 5) is executed, for the first arriving item $e$, we have $v_{i}(e)=a_{i}$.
@@
The reason why the PBC algorithm is executed is that for the arriving item $e$, if we add it to agent $i$’s bundle, there will be an envy cycle between two agents. i.e., $v_{i}(A_{i}\cup\{e\})<v_{i}(A_{j})$ and $v_{j}(A_{j})<v_{i}(A_{i}\cup\{e\})$. Suppose for the contradiction that $v_{i}(e)=b_{i}$. Since the input allocation $A$ is EF1, for agent $i$, we have $v_{i}(A_{j})-v_{i}(A_{i})\leq b_{i}$. Thus, allocating item $e$ to agent $i$ eliminates the envy and cannot lead to an envy cycle, which is a contradiction. Therefore, for the arriving item $e$, it holds that $v_{i}(e)=a_{i}$. ∎
@@
When the PBC algorithm (Algorithm 5) is executed, we have $|A_{i}|\geq 1$ and $|A_{j}|\geq 1$.
@@
Suppose that $|A_{i}|=0$. In that case, we have $|A_{j}|=1$. Otherwise, there would be an envy cycle. After allocating $e$, there will be an envy cycle, i.e., $v_{i}(e)=a_{i}<v_{i}(A_{j})=b_{i}$ and $v_{j}(A_{j})<v_{j}(e)$. If the second inequality holds, we have $v_{j}(A_{j})=a_{j}$ but $v_{i}(A_{j})=b_{i}$, which contradicts the line 4 in the EGP algorithm (Algorithm 4). Suppose that $|A_{j}|=0$. In that case, we have $|A_{i}|=0$, which contradicts the fact that after allocating $e$, an envy cycle appears. ∎
@@
The PBC algorithm (Algorithm 5) computes an NW allocation satisfying:
  • $\frac{1}{2}$-EF1,
  • or EF1 with an envy cycle,
  • or EF1 without an envy cycle.
@@
When the PBC algorithm is implemented ($\alpha=2$), the input allocation $\textbf{A}=(A_{1},A_{2})$ output by the EGP algorithm (Algorithm 4) is EF1 without an envy cycle. By the description of the PBC algorithm, it is executed in at most two consecutive rounds.
@@
$\lambda_{1}=1$. For the first arriving item $e_{1}$, by Observation 1, we have $v_{i}(e_{1})=a_{i}$. If $\lvert A_{j}\rvert=1$, we have $v_{i}(A_{j})=b_{i}$ since agent $i$’s bundle is EF1. Then, agent $j$ picks this item. For agent $i$, her bundle is EF1 since we have $\lvert A_{i}\lvert\geq 1$ and $v_{i}(e_{1})=a_{i}$. For agent $j$, her new bundle is EF since she does not envy agent $i$. Thus, this allocation is EF1 without an envy cycle, and the algorithm is terminated. If $\lvert A_{j}\rvert>1$, let us consider two cases about the valuation of item $e_{1}$.
@@
  • If $v_{j}(e_{1})=b_{j}$, item $e$ is allocated to agent $j$. Then, we set $\lambda_{1}=2$ and $\mu_{1}=1$. In this case, for agent $j$, her new bundle is EF since $v_{j}(A_{j}\cup\{e_{1}\})>v_{j}(A_{j})\geq v_{j}(A_{i})$ and for agent $i$, her bundle is $\frac{1}{2}$-EF1, i.e, $v_{i}(A_{i})\geq\frac{1}{2}v_{i}(A_{j}\cup\{e_{1}\}\setminus\{g\})$ holds for $g\in A_{j}\cup\{e_{1}\}$, since we have $v_{i}(A_{i})\geq v_{i}(A_{j}\setminus\{g\})$ for some $g\in A_{j}$ and $v_{i}(A_{i})\geq a_{i}=v_{i}(e_{1})$. Thus, this allocation is $\frac{1}{2}$-EF1.
@@
  • If $v_{j}(e_{1})=a_{j}$, item $e$ is allocated to agent $i$. Then, we set $\lambda_{1}=2$ and $\mu_{1}=2$. In this case, for agent $i$, her new bundle is still EF1 since she is an unenvied agent in Algorithm 3 and the input allocation $A$ is EF1, and for agent $j$, her bundle is EF1 since we have $v_{j}(A_{j})\geq v_{j}(A_{i})$ but $v_{j}(A_{j})<v_{j}(A_{j}\cup\{e_{1}\})$. Thus, this allocation is EF1 with an envy cycle.
It is clear that any possible allocation from the above cases does not satisfy EF1 without an envy cycle. Thus, we keep $\alpha=2$.
$\lambda_{1}=2$. The second item $e_{2}$ arrives. Let us consider two cases about $\mu_{1}$.
@@
  • If $\mu_{1}=1$, agent $i$ picks it. In this case, for agent $j$, her bundle is still EF since we have $v_{j}(A_{j})\geq v_{j}(A_{i})$ and $v_{j}(e_{1})=b_{j}\geq v_{j}(e_{2})$. For agent $i$, her new bundle is EF1 since we have $v_{i}(A_{i})\geq v_{i}(A_{j}\setminus\{g\})$ for some $g\in A_{j}$ and $v_{i}(e_{2})\geq v_{i}(e_{1})=a_{i}$. Thus, this allocation is EF1 without an envy cycle, and we set $\alpha=1$, $\lambda_{1}=1$, and $\mu_{1}=0$.
@@
  • If $\mu_{1}=2$, let us consider two cases. When $v_{i}(e_{2})=a_{i}$, agent $j$ picks it. For agent $i$, her bundle is still EF1 since we have $v_{i}(A_{i})\geq v_{i}(A_{j}\setminus\{g\})$ for some $g\in A_{j}$ and $v_{i}(e_{1})=v_{i}(e_{2})=a_{i}$. For agent $j$, her new bundle is EF because $v_{j}(A_{j})\geq v_{j}(A_{i})$ and $v_{j}(e_{2})=v_{j}(e_{1})=a_{j}$. Hence, the allocation is EF1 without an envy cycle, and we set $\alpha=1$, $\lambda_{1}=1$, and $\mu_{1}=0$. When $v_{i}(e_{2})=b_{i}$, it is allocated to agent $i$. For agent $i$, her new bundle is EF since it holds that $v_{i}(A_{i}\cup\{e_{1},e_{2}\})=v_{i}(A_{i})+a_{i}+b_{i}\geq v_{i}(A_{j}\setminus\{g\})+v_{i}(g)+a_{i}$ for some $g\in A_{j}$ where the inequality follows from the fact that $v_{i}(g)\in\{a_{i},b_{i}\}$. For agent $j$, her bundle is $\frac{1}{2}$-EF1, i.e., $v_{j}(A_{j})\geq\frac{1}{2}v_{j}(A_{i}\cup\{e_{1},e_{2}\}\setminus\{g\})$ holds for some $e\in A_{i}\cup\{e_{1},e_{2}\}$, since $v_{j}(A_{j})\geq v_{j}(A_{i})$ and $v_{j}(A_{j})\geq v_{j}(e_{1})=\min\{v_{j}(e_{1}),v_{j}(e_{2})\}$. Thus, this allocation is $\frac{1}{2}$-EF1, and we set $\alpha=3$, $\lambda_{1}=1$, and $\mu_{1}=0$.
@@
When the DBC algorithm (Algorithm 6) is executed, there are two properties about input allocation $\bf A$:
@@
  • Let $e_{1}$ and $e_{2}$ denote two new allocated items in the PBC algorithm, we have $v_{j}(A_{j})\geq\frac{1}{2}v_{j}(A_{i}\setminus\{e\})$, where $e\in\arg\max_{g\in\{e_{1},e_{2}\}}v_{j}(g)$.
@@
When the DBC algorithm is ready to be executed, the PBC algorithm is terminated for $\lambda_{1}=\mu_{1}=2$. Then, by the proof of Lemma 5, it is easy to see that the statements are correct. ∎
@@
The DBC algorithm (Algorithm 6) computes an NW allocation satisfying
  • $\frac{1}{2}$-EF1,
  • or EF1 with an envy cycle,
  • or EF1 without an envy cycle.
@@
By Observation 3, the input allocation $\textbf{A}=(A_{1},A_{2})$ is $\frac{1}{2}$-EF1, where $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}$ and $v_{j}(A_{j})\geq\frac{1}{2}v_{j}(A_{i}\setminus\{e\})$ for some $e\in A_{i}$. Unlike the PBC algorithm, this algorithm may be executed in many rounds. Let $g_{1}$ and $g_{2}$ denote the two online items that are allocated to agent $i$ in the PBC algorithm and we have $\min\{v_{j}(g_{1}),v_{j}(g_{2})\}=a_{j}$. Currently, we have $\alpha=3$.
$\lambda_{2}=1$. For the arriving item $e_{1}$ when $\lambda_{2}=1$, let us consider three cases of agents’ valuations.
  • When $v_{i}(e_{1})=a_{i}$, agent $j$ picks it. There are two subcases about $v_{j}(e_{1})$. If $v_{j}(e_{1})=b_{j}$, for agent $i$, her bundle is EF since we have $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}=v_{i}(A_{j}\cup\{e_{1}\})$. For agent $j$, her new bundle is EF1 since we have $v_{j}(A_{j})\geq v_{j}(A_{i}\setminus\{g_{1},g_{2}\})$ and $v_{j}(e_{1})=b_{j}\geq\max\{v_{j}(g_{1}),v_{j}(g_{2})\}$. Thus, this allocation is EF1 without an envy cycle. If $v_{j}(e_{1})=a_{j}$, for agent $i$, her bundle is EF since we have $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}=v_{i}(A_{j}\cup\{e_{1}\})$. For agent $j$, her new bundle is EF1 since we have $v_{j}(A_{j})\geq v_{j}(A_{i}\setminus\{g_{1},g_{2}\})$ and $v_{j}(e_{1})=a_{i}=\min\{v_{j}(g_{1}),v_{j}(g_{2})\}$. Thus, this allocation is EF1 without an envy cycle. Combining the above two cases, we set $\alpha=1$, $\lambda_{2}=1$, and $\mu_{2}=1$, implying that the EGP algorithm is ready to be executed.
  • When $v_{i}(e_{1})=b_{i}$ and $v_{j}(e_{1})=b_{j}$, it is allocated to agent $j$ and we set $\mu_{2}=1$. For agent $i$, her bundle is EF1 since after removing $e_{1}$ from $A_{j}$, we have $v_{i}(A_{i})\geq v_{i}(A_{j})+a_{i}$. For agent $j$, her new bundle is also EF1 since we have $v_{j}(A_{j})\geq v_{j}(A_{i}\setminus\{g_{1},g_{2}\})$ and $v_{j}(e_{1})=b_{j}\geq\max\{v_{j}(g_{1}),v_{j}(g_{2})\}$. Thus, this allocation is EF1 with an envy cycle, and we set $\lambda_{2}=2$.